A 1.6-cm-thick wood floor covers a 4.0m×5.5m room.

The subfloor on which the flooring sits is at a temperature of 16.2 celsius degrees, while the air in the room is at 19.1 celsius degrees. What is the rate of heat conduction through the floor? Use 0.200 W/m⋅K for the thermal conductivity of wood.

Express your answer using two significant figures.
Answer #1

given

length is L = 1.6 cm

= 0.016 m

the Area is A = 4.0m x 5.5m

= 22 m2

change in temperature is dT = 19.1 - 16.2

= 2.9 K

thermal conductivity of wood, k = 0.200 W/m.K

using equation for rate of heat conduction ( Q/t )

Q/t = k A dT / L

= 0.2 x 22 x 2.9 / 0.016

Q/t = 797.50 J/sec

the rate of heat conduction through the floor is Q/t = 797.50 J/sec